OpenStudy (anonymous):

Algebra 2. Fan and Medal

OpenStudy (anonymous):

question?

OpenStudy (anonymous):

lol you want us to do it for you

OpenStudy (anonymous):

why not?

OpenStudy (anonymous):

No. You dont have to. I just didnt understand it and I need help.

OpenStudy (anonymous):

to find the equation of the line between the points \((0,3)\) and \((1,2)\) you need the slope

OpenStudy (anonymous):

so i would use y=mx+b?

OpenStudy (anonymous):

right one, down one means the slope is \(-1\) since it goes through the point \((0,3)\) you know the \(y\) intercept is \(3\) making your equation \[y=-x+3\]

OpenStudy (anonymous):

yes, \(y=mx+b\) with \(m=-1,b=3\)

OpenStudy (anonymous):

oh damn they said "standard form" hold the phone

OpenStudy (anonymous):

so it would look like y=(-1)x+3?

OpenStudy (anonymous):

\[y=-x+3\\ x+y=3\] second one is standard form

OpenStudy (anonymous):

yes, it is \(y=(-1)x+3\) but if you write it that way, displaying the \(-1\) your teacher will think you are an idiot, so don't do it, just write \(y=-x+3\) then change it to \(x+y=3\)

OpenStudy (anonymous):

lol alright. Sorry I suck at math D:

OpenStudy (anonymous):

yeah its okay just don't put a 1 or a -1 in front of anything wanna do the second one?

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

points are \((0,3)\) and \((-3,0)\) slope is \[\frac{3-0}{0-(-3)}=\frac{3}{3}=1\]

OpenStudy (anonymous):

again the point \((0,3)\) tells you the \(y\) intercept is \(3\) so write \[y=mx+b\] with \(m=1,b=3\)

OpenStudy (anonymous):

but whatever you do, do not write \[y=\color{red}1x+3\] just write \[y=x+3\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw might as well finish them right? \[(-3,0),(-1,-4)\] slope is \[\frac{0-(-4)}{-3-(-1)}=\frac{4}{-2}=-2\]

OpenStudy (anonymous):

this time you do not know the \(y\) intercept so you have to use the point - slope formula \[y-y_1=m(x-x_1)\] with \[x_1=-3,y_1=0, m=-2\]

OpenStudy (anonymous):

y-0=-2(x--3) ?

OpenStudy (anonymous):

yes, now lets clean it up

OpenStudy (anonymous):

\[y=-2(x+3)\] would be the first step then distribute the \(-2\) to finish

OpenStudy (anonymous):

you should end up with \[y=-2x-6\] hope this was more or less clear finding the equation of a line given two points is not too hard and a decent skill to have

OpenStudy (anonymous):

You helped me alot. Gave me a better understanding. Thank you so much :)

OpenStudy (anonymous):

yw